C97TA

PI: Sanghera
Title: 0016+731, mode A test

Test: Jan 1997
Frequency: 6 cm
Mode: A
Obs time (UT): 13 EU 15:15(13/01)-15:45(13/01)
Stations: EbJbMcOnWb
Processor: Bonn
EU Mk3 Tapes/St.: 1

Comment


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (dgraham at Mon Jan 13 10:25:26 EST 1997) Minor failures
system time was 5 minutes different to cpu. started 5 minutes late. So both scans will be short. Problems with phase cal.

Westerbork (tony foley at Mon Jan 13 11:39:22 EST 1997) Success
Used telescope RT7 at position: X=3828647.18 Y=-443450.02 Z=5064923.08 (this is ~144m east of RT6, used in December) System temp equivalent ~ 2400Jy Linear polarization NO phasecal (unit is broken) Note also that VC11 and VC12 have been swapped from their positions in the December tests.

Jodrell (Alastair Gunn at Tue Jan 14 05:38:32 EST 1997) Success
No known problems. Apparently everything worked okay. Weather overcast, dry, occasional sunny periods, wind calm. Tape JODR0138,DB41 shipped to Bonn with Woodcroft on Tuesday 14th 1997.

Medicina (Tiziana Venturi at Tue Jan 14 10:32:28 EST 1997) Minor failures
The experiment ran smoothly. Apparently, no problems occurred.

Onsala (John Conway at Tue Jan 14 11:27:41 EST 1997) Success
Few minutss lost because of a stuck head on second pass. No other known problems